A jeweler paid 1455 for a watch. He wants to price the watch for sale so that he can give the customer a 25% discount on the selling price and still make a profit of 30% on his cost. What selling price should he mark for the watch? Answer by checkley71(8403) 
 1455*1.3=1891.5 IS THE PRICE FOR A 30% PROFIT. THEN
1891.5=(1891.5+X).75 OR 1891.5=1418.625+.75X OR 472.875=.75X OR X=472.875/.75
OR X=630.5 THE THE SELLING PRICE WITH A 25% DISCOUNT THAT GIVES A 30% PROFIT
IS 1891.5+630.5 OR $2522=SELLING PRICE BEFORE DISCOUNT.